# Real-World Data and Calibrated Simulation Suite for Offline Training of Reinforcement Learning Agents to Optimize Energy and Emission in Buildings for Environmental Sustainability

### v1

@@ B.4 Reward Data Protos
  • The ZoneRewardInfo message provides information about the zone air temperature measurements, temperature setpoints, airflow rate and setpoint, and average occupancy for the time step. Each instance is indexed by its unique zone ID.
  • The AirHandlerRewardInfo message describes the combined electrical power in W use of the intake/exhaust blowers, and the electrical power in W of the refrigeration cycle. Since a building may have more than one air handler, the air handler objects are values in a map keyed by the air handlers’ device IDs.
@@
B.4.1 RewardInfo Definition
1message RewardInfo {
2 // Information about each zone in the time step for computing reward.
3 message ZoneRewardInfo {
4 // Heating setpoint of the zone at the timestep in K.
5 float heating_setpoint_temperature = 1;
6
7
8 // Cooling setpoint of the zone at the timestep in K.
9 float cooling_setpoint_temperature = 2;
10
11
12 // Average zone air temperature measured in the zone in K.
13 float zone_air_temperature = 3;
14
15
16 // Setpoint for air flow ventilation in the zone in m^3/s.
17 float air_flow_rate_setpoint = 4;
18
19
20 // Actual ventilation air flow in the zone in m^3/s.
21 float air_flow_rate = 5;
22
23
24 // Average occupancy in the zone over the time step in number of
25 // people in the zone.
26 float average_occupancy = 6;
27 }
28
29 // Information about the air handler energy consumption for computing reward.
30 message AirHandlerRewardInfo {
31 // Cumulative electrical power in W applied to blowers.
32 float blower_electrical_energy_rate = 1;
33
34 // Cumulative electrical energy rate applied in W for air conditioning. This
35 // represents the total power applied for running a refrigeration or
36 // heat pump cycles (includes running a compressor and pumps to
37 // recirculate refrigerant.).
38 float air_conditioning_electrical_energy_rate = 2;
39 }
40
41
42 // Information about the boiler that provides heated water for VAVs.
43 message BoilerRewardInfo {
44 // Energy rate consumed in W by natural gas for heating water.
45 float natural_gas_heating_energy_rate = 1;
46 // Cumulative electrical power in W for water recirculation pumps.
47 float pump_electrical_energy_rate = 2;
48 }
49
50
51 // Start and end timestamps bound the timestep of the reward information.
52 google.protobuf.Timestamp start_timestamp = 1;
53 google.protobuf.Timestamp end_timestamp = 2;
54
55
56 // Unique ID of the agent (controller). This should reflect the
57 // attributes of the RL models, including the type of algo and its
58 // parameters.
59 string agent_id = 3;
60
61
62 // Unique ID of the scenario being executed. This should reflect the details
63 // of the scenario. In simulation, it should identify the canonical scenario.
64 // In real world, it should define the building and start date/time.
65 string scenario_id = 4;
66
67
68 // Map with zone_id and zone reward info for all zones in the building
69 // under control of the agent. The zone_id could be a unique room number,
70 // or the specific zone coordinates: (i.e., 'z_i,z_j') from the simulation.
71 map<string, ZoneRewardInfo> zone_reward_infos = 5;
72
73
74 // Information about the air handlers' energy consumption required to
75 // calculate the reward.
76 map<string, AirHandlerRewardInfo> air_handler_reward_infos = 6;
77
78
79 // Information about the boilers' energy consumption required to compute the
80 // reward.
81 map<string, BoilerRewardInfo> boiler_reward_infos = 7;
82}
B.4.2 RewardInfo Example
1start_timestamp {
2 seconds: 1681109700
3}
4end_timestamp {
5 seconds: 1681110000
6}
7agent_id: "baseline_policy"
8scenario_id: "baseline_collect"
9zone_reward_infos {
10 key: "rooms/1000004614278"
11 value {
12 heating_setpoint_temperature: 289.0
13 cooling_setpoint_temperature: 298.0
14 zone_air_temperature: 293.5944519042969
15 air_flow_rate_setpoint: 258.0
16 air_flow_rate: 12.0
17 }
18}
19zone_reward_infos {
20 key: "rooms/1000004658174"
21 value {
22 heating_setpoint_temperature: 289.0
23 cooling_setpoint_temperature: 298.0
24 zone_air_temperature: 293.4277648925781
25 air_flow_rate_setpoint: 60.0
26 }
27}
28zone_reward_infos {
29 key: "rooms/1000004658175"
30 value {
31 heating_setpoint_temperature: 289.0
32 cooling_setpoint_temperature: 298.0
33 zone_air_temperature: 293.03887939453125
34 air_flow_rate_setpoint: 185.0
35 air_flow_rate: 4.001242637634277
36 }
37}
38zone_reward_infos {
39 key: "rooms/1000004658176"
40 value {
41 heating_setpoint_temperature: 289.0
42 cooling_setpoint_temperature: 298.0
43 zone_air_temperature: 293.53887939453125
44 air_flow_rate_setpoint: 145.0
45 air_flow_rate: 53.0
46 }
47}
48air_handler_reward_infos {
49 key: "12945159110931775488"
50 value {
51 }
52}
53air_handler_reward_infos {
54 key: "14409954889734029312"
55 value {
56 }
57}
58boiler_reward_infos {
59 key: "13761436543392677888"
60 value {
61 pump_electrical_energy_rate: 1527.1470947265625
62 }
63}
B.4.3 RewardResponse Definition
1// The return reward signal from the reward function. While the principal
2// signal is the agent reward and should be returned to the RL agent, the
3// other fields provide useful information for tracking and monitoring.
4// One EnergyRewardResponse is associated with each EnergyRewardInfo.
5message RewardResponse {
6
7
8 // Complete reward signal to be returned to the agent.
9 float agent_reward_value = 1;
10
11
12 // Cumulative productivity is measured in USD, and represents the total
13 // estimated productivity of the building.
14 float productivity_reward = 2;
15
16
17 // Total electrical energy cost estimate in USD.
18 float electricity_energy_cost = 3;
19
20
21 // Total natural gas energy cost in USD.
22 float natural_gas_energy_cost = 4;
23
24
25 // Estimated carbon emitted in kg.
26 float carbon_emitted = 5;
27
28
29 // Estimated carbon cost in USD.
30 float carbon_cost = 6;
31
32
33 // Productivity weight parameter.
34 float productivity_weight = 7;
35
36
37 // Energy Cost Weight parameter.
38 float energy_cost_weight = 8;
39
40
41 // Carbon emission weight parameter.
42 float carbon_emission_weight = 9;
43
44
45 // Productivity factor (avg labor value of one person-hour).
46 float person_productivity = 10;
47
48
49 // Total average occupancy across all zones.
50 float total_occupancy = 11;
51
52
53 // Reward scale for normalizing the reward
54 float reward_scale = 12;
55
56
57 // Reward shift for normalizing the reward
58 float reward_shift = 13;
59
60
61 // Total productivity regret = max productivity - actual productivity
62 float productivity_regret = 14;
63
64
65 // Normalized productivity regret
66 float normalized_productivity_regret = 15;
67
68
69 // Normalized energy cost =
70 // combined_energy_cost /
71 // (max_electricity_energy_cost + max_natural_gas_energy_cost)
72 float normalized_energy_cost = 16;
73
74
75 // Normalized carbon emission =
76 // combined_carbon_emission /
77 // (max_electricity_carbon_emission + max_natural_gas_carbon_emission)
78 float normalized_carbon_emission = 17;
79
80
81 // Start and end timestamps bound the timestep of the reward information.
82 google.protobuf.Timestamp start_timestamp = 18;
83 google.protobuf.Timestamp end_timestamp = 19;
84}
B.4.4 RewardResponse Example
1agent_reward_value: -0.00222194055095315
2electricity_energy_cost: 0.022907206788659096
3carbon_emitted: 0.011416268534958363
4productivity_weight: 0.5
5energy_cost_weight: 0.20000000298023224
6carbon_emission_weight: 0.30000001192092896
7person_productivity: 300.0
8reward_scale: 1.0
9normalized_energy_cost: 0.0090464623644948
10normalized_carbon_emission: 0.0013754934770986438
11start_timestamp {
12 seconds: 1681109700
13}
14end_timestamp {
15 seconds: 16811100
Appendix C Simulator Design Consideration Details
A simulator models the physical system dynamics of the building, devices, and external weather conditions, and can train the control agent interactively, if the following desiderata are achieved:
@@ Appendix C Simulator Design Consideration Details
Appendix D Derivation for Tensorized Finite Difference (FD) Equations
@@
D.1 Assembling the Energy Balance
@@
D.2 Computing heat transfer via conduction, convection, and thermal absorption
@@
The negative sign in Equations 4 - 9 are due to the fact that energy flows in the direction opposite of the temperature gradient, $\Delta T$, i.e., from high to low.
@@
    Exterior CVs are CVs that represent the ambient weather conditions, such as $T_{\infty}$ , which are note calculated by the FD calculator, just specified by the current input conditions.
@@
D.3 Solving for the temperature at each CV
@@
Combining the Energy Balance in Equation 2 with the conduction and convection equations (Equations 7-10) we can include all terms for all faces on the $i,j$ CV. Our goal is to solve for $T_{i,j}$ which can then be run over multiple sweeps to convergence.
@@
D.4 Tensorizing the temperature estimate
@@
For each timestep, we execute Equation D.4 as single-step tensor operations until convergence, where the maximum change across all CVs between current and last iteration is less then a conservative lower threshold, $\epsilon\leq 0.01^{\circ}C$
Appendix E Simulator Configuration Procedure Details
@@
    HVAC metadata. This includes each device, its name, location, setpoints, fixed parameters and purpose.
We preprocess the detailed floorplan blueprints of the building, and extract a grid that gives us an approximate placement of walls and how rooms are divided. This is done via the following procedure:
@@
We tested this pipeline on SB1, which consisted of two floors with combined surface area of 93,858 square feet, and has 127 HVAC devices. Given floorplans and HVAC layout information, a single technician was able to generate a fully specified simulation in under three hours. This customized simulator matched the real building in every device, room, and structure.
Appendix F Calibration Hyperparameter Tuning Details
@@
Appendix G Additional Spatial Error Visualizations
@@
Appendix H Simulator SAC Agent Training Details and Performance Analysis
@@
Appendix I Training and Evaluating a Learned Dynamics Model
Aside from being useful for offline training and for calibrating our simulator, the real world data can also be used to directly learn a regression model that approximates the building dynamics. This model can then be used to train a control agent.
As described in the main paper, to demonstrate this approach, building off of earlier work(Velswamy et al., 2017; Sendra-Arranz & Gutiérrez, 2020; Zou et al., 2020; Zhuang et al., 2023), we trained an LSTM to model the building dynamics. We used an encoder-decoder network, where the model takes in a historical sequence of length $N$ and outputs a prediction sequence of length $M$. At each timestep $t$ in the sequence, the model is given an observation $O_{t}$, action taken by the policy $A_{t}$, and auxiliary state features (such as time of day and weather, that are useful as inputs but need not be predicted) $U_{t}$, and for future timesteps, the model is trained to predict future observations, as well as future reward information (based on predicted energy use and carbon emissions) $E_{t}$. The LSTM model is shown in Figure 24.
Figure 24: Architecture of LSTM building dynamics model
We then trained the model to predict the next observation for 65 epochs, plotting training and validation loss, as shown in figure 25
Figure 25: Loss of LSTM building dynamics model, with train loss in orange and validation loss in blue.
However, loss curves alone do not tell the full story of how well our regression model is reconstructing the signal of the dynamics, so we also included additional evaluations. We had the model predict 3 weeks into the future, and then compared the predictions with the ground truth data to ensure the cyclic patterns of the medians are reproduced. The chart in figure 26 shows 20 measurement time series from the regression models shown in yellow compared to the actual values shown in gray. By inspection, we conclude that the regression building provides good correspondence with the actual real data signals.
Figure 26: Detailed analysis of learned dynamics as compared to real data.
Appendix J Real Data SAC Agent Training Details and Performance Analysis
We then trained a SAC agent on the regression environment, much like how we did on the simulator. This gives us a baseline for how to generate a policy purely based on data, without use of the simulator. We used hyper-parameter tuning, and trained 200 agents. The chart in figure 27 shows agent reward progress as the number of trials increased.
Figure 27: Detailed analysis of learned dynamics as compared to real data.
To compare the learned policy with the baseline, we plotted the two policies in 28. The baseline and agent episode temperature timelines shown below provide a temporal perspective of the environment median zone air temperatures (yellow) and setpoints (white), outside air temperature (blue), and the agent actions on the environment (water temperature setpoints (lime), and air handler temperature setpoints (magenta). While the regression model under baseline policy correctly represents the weekend setpoint ranges, the regression building applies nearly the weekday setpoint ranges when running under agent control. This is likely due to the agent applying setpoints that regression associates with weekday actions, and incorrectly returns a setpoint that is closer to the weekday. For this reason, we do not evaluate the model’s performance on weekends. Similar to baseline control, the agent ramps up water temperature (lime) at the beginning of the day. However, the agent tends to maintain the water temperature around 80C for substantially longer than baseline control. At first glance, this may seem counterproductive. However, heat exchange is also based on water flow and air flow. Lower supply water temperatures require more airflow to transfer the same amount of heat. Therefore, higher water temperatures do not necessarily result in higher energy consumption. Also, note that the agent does not drop the water temperature as low as the baseline policy, and the agent tends to apply smoother actions compared to the baseline’s rapid oscillation between 40 and 60C. We speculate that one strength of the proposed solution is the agent’s ability to discover better and non-intuitive policies that are unlikely to be chosen by human HVAC technicians. The agent also has a different control policy for the air handlers’ supply air temperatures, shown in magenta. On one air handler’s supply air temperature, the agent tends to operate SB1:AHU:AC 1 at a higher temperature than SB1:AHU:AC 2.
Figure 28: Detailed analysis of learned dynamics as compared to real data.
Finally, much like how we did with the simulated agent, we break down the reward into its four components and see how the agent did relative to the baseline on the regression building model.
J.1 Setpoint Management Performance
The difference in setpoint deviation between agent and baseline was insignificant. However, at 23C the average setpoint deviation was slightly higher, but was still within a narrow window (less than 1/10 C). The setpoint deviation test using the regression model may be slightly optimistic compared to the real building, because the regression model only approximates the zone temperatures with a single median, hiding the larger spread of temperatures throughout the building.
Figure 29: Setpoint Management Performance.
J.2 Carbon Emission Performance
In 12 of 19 temperature bins the agent generated less carbon than the baseline. While only two temperature bins (17C, 25C) resulted in confidence greater than 90%, the results indicate a reduction in carbon emission on most of the bins. The agent tends to emit more carbon in the moderate temperature ranges (21, 22C), likely due to a higher setpoint during the day than the baseline. Overall, the agent performs favorably, even though most bins have a low statistical confidence.
Figure 30: Carbon Emission Performance.
J.3 Electrical Energy Performance
While no temperature bins yielded confidence scores greater than 90%, the agent tends to consume less electricity than the baseline, except for the 21, 22C temperature bins. Under both policies, electricity consumption dramatically increases with outside air temperature.
Figure 31: Electrical Energy Performance.
J.4 Natural Gas Energy Performance
The agent policy tends to consume less natural gas than the baseline policy, even though only three yielded significant reduction with confidence of at least 90% (17, 24, 25C).
Figure 32: Natural Gas Energy Performance.

### v2

@@ C.6 Reward Data
  • The ZoneRewardInfo provides information about the zone air temperature measurements, temperature setpoints, airflow rate and setpoint, and average occupancy for the timestep. Each instance is indexed by its unique zone ID.
  • The AirHandlerRewardInfo describes the combined electrical power in W use of the intake/exhaust blowers, and the electrical power in W of the refrigeration cycle. Since a building may have more than one air handler, the air handler objects are values in a map keyed by the air handlers’ device IDs.
@@
The data is again stored as matrices, one of RewardInfo and one of RewardResponse. The RewardInfo matrix is dimension TxR, where T is the number of timesteps and R the number of factors used to calculate the reward, again with a list of T timesteps and R ids that determine which device and meter each column refers to. Similarly, the RewardResponse is Dimension TxC, where T is timesteps as before, and C is the number of reward function constants + 1 for the scalar reward itself.
Appendix D Simulator Design Considerations
Every building is unique, due to its physical layout, equipment, and location. Fully customizing a high-fidelity simulation to a specific target building requires nearly exhaustive knowledge of the building structure, materials, location, etc., some of which are unknowable, especially for legacy office buildings. This requires manual “guesstimation” which can erode the accuracy promised by high-fidelity simulation. In general, the configuration time required for high-fidelity simulations limits their utility for deploying RL-based optimization to many buildings. High-fidelity simulations also are affected by computational demand and long execution times.
Alternatively, we propose a fast, low-to-medium-fidelity simulation model that was useful in addressing various design decisions, such as the reward function, the modeling of different algorithms. and for end-to-end testing. The simulation is built on a 2D finite-difference (FD) grid that models thermal diffusion, and a simplified HVAC model that generates or removes heat on special “diffuser” CV in the FD grid. While the uncalibrated simulator is of low-to-medium fidelity, the key additional factor is data. We collect recorded observations from the target building under baseline control, and use that data to calibrate the simulator, by adjusting the simulator’s physical parameters to minimize difference between real and simulated data. We believe this approach hits the sweet spot in this tradeoff, enabling scalability while maintaining a high enough level of fidelity to train an improved policy.
Figure 9: Simulation Fidelity vs. Execution Speed. The ideal operating point for training RL agents for energy and emission efficiency is a tradeoff between fidelity, depicted as 1 minus a normalized error $\epsilon$ between simulation and real, and execution speed, as measured by the number of training steps per second. Additional consideration also includes the time to configure a custom simulator for the target building. While many approaches tend to favor high fidelity over execution, speed, our approach argues a low-to-medium fidelity that has a medium-to-high speed is most suitable for training an RL agent.
Thus, a simulator models the physical system dynamics of the building, devices, and external weather conditions, and can train the control agent interactively, if the following desiderata are achieved:
@@ Appendix D Simulator Design Considerations
Appendix E Derivation for Tensorized Finite Difference (FD) Equations
@@
E.1 Assembling the Energy Balance
@@
E.2 Computing heat transfer via conduction, convection, and thermal absorption
@@
The negative sign in Equations 7 - 12 are due to the fact that energy flows in the direction opposite of the temperature gradient, $\Delta T$, i.e., from high to low.
@@
    Exterior CVs are CVs that represent the outside weather conditions, such as $T_{\infty}$ , which are note calculated by the FD calculator, just specified by the current input conditions.
@@
E.3 Solving for the temperature at each CV
@@
Combining the Energy Balance in Equation 5 with the conduction and convection equations (Equations 10-13) we can include all terms for all faces on the $i,j$ CV. Our goal is to solve for $T_{i,j}$ which can then be run over multiple sweeps to convergence.
@@
E.4 Tensorizing the temperature estimate
@@
For each timestep, we execute Equation E.4 as single-step tensor operations until convergence, where the maximum change across all CVs between current and last iteration is less then a conservative lower threshold, $\epsilon\leq 0.01^{\circ}C$
Appendix F Simulator Configuration Procedure Details
@@
    HVAC metadata. This includes each device, its name, location, setpoints, fixed parameters, and purpose.
We preprocess the detailed floorplan blueprints of the building and extract a grid that gives us an approximate placement of walls and how rooms are divided. This is done via the following procedure:
@@
We tested this pipeline on SB1, which consisted of two floors with combined surface area of 93,858 square feet, and has 173 HVAC devices. Given floorplans and HVAC layout information, a single technician was able to generate a fully specified simulation in under three hours. This customized simulator matched the real building in every device, room, and structure.
Appendix G Calibration Hyperparameter Tuning Details
@@
Appendix H Additional Spatial Error Visualizations
@@
Appendix I Additional ModNN Details
I.1 Detailed ModNN Model Structure
In the ModNN model, a Gated Recurrent Unit (GRU) module is used to capture the highly nonlinear relationships between disturbance variable inputs and heat gains, while a Fully Connected (FC) neural network module captures the HVAC input. Notably, the HVAC input in this study is air-side, allowing it to be directly added to the disturbance variables. However, this module can be extended to accommodate different HVAC systems. For instance, a radiation-based HVAC system would require a Recurrent Neural Network (RNN) module to consider the heat lagging effects.
Figure 20: Diagram of Detailed ModNN Model Structure.
The integrated heat transfer terms will go through another FC module to model the dynamics of zone air mass and heat capacity. The output represents the temperature change per timestep, which is recursively added to the previous timestep’s temperature for future predictions.
I.2 Adjacent Matrix for Multizone Dynamic Modeling
To extend the single-zone model to a multi-zone framework, we developed a conduction heat transfer model to capture interactions between adjacent zones. A skew-symmetric matrix is used to represent temperature differences between adjacent zones, as shown in the figure. Based on the heat conduction equation:
$Q=\frac{kA\Delta(T_{i}-T_{j})}{l}:=f_{NN_{adj}}(T_{i}-T_{j})$
Where $q$ represents heat transfer through conduction, $k$ is the thermal conductivity of the material, $A$ is the cross-sectional area, $T_{i}$ and $T_{j}$ are the space air temperatures of the adjacent zones, and $l$ is the thickness of the wall.
Figure 21: Adjacency Matrix for Extending from Single-Zone to Multi-Zone.
This heat transfer term is learned using another FC layer and then integrated with the other heat transfer terms to predict the next step’s space air temperature.
I.3 Data Structure
The input dimension of the encoder is
$(D_{\text{state}}+D_{\text{dis}}+D_{\text{adj}}+D_{\text{hvac}})\times B\times L_{\text{en}},$
where $D_{\text{state}}$, $D_{\text{dis}}$, $D_{\text{adj}}$, and $D_{\text{hvac}}$ are the feature dimensions of the state variables (the number of zones), disturbance variable, number of adjacent zones, and control variables, respectively. $B$ is the batch size, and $L_{\text{en}}$ is the length of the encoder.
The input dimension of the decoder is
$(D_{\text{dis}}+D_{\text{adj}}+D_{\text{hvac}})\times B\times L_{\text{de}},$
since the future indoor air temperature is unknown at timestep $t$. Thus, the input dimension for the decoder excludes $D_{\text{state}}$.
The output dimension is
$1\times B\times L_{\text{de}}.$
The input vector can be formulated as shown below:
Here, we define $w^{k}_{t}\in\mathbb{R}^{\text{Disturbance}\times{\text{batch}\times(L_{\text{en}}+L_{\text{de}})\times D}}$ as the disturbance vector, where $t$ is the timestep and $k$ is the feature category. The control input term is represented as
$u^{k}_{t}\in\mathbb{R}^{1\times\text{batch}(L_{\text{en}}+L_{\text{de}})\times 1}$
and the state variable, representing space air temperature in building dynamics modeling, is given by
$x^{k}_{t}\in\mathbb{R}^{1\times\text{batch}\times(L_{\text{en}}+1)\times 1}.$
We further clarify the abbreviations for $k$ used in this matrix:
- $solar$: global solar radiation
- $time$: time information
- $T_{\text{OA}}$: outside air temperature
- $Occ$: occupant number
I.4 Hyperparameters
We summarized all the hyperparameters used in this study as shown below. The input vector can be formulated as shown below:
Appendix J Simulator SAC Agent Training Details and Performance Analysis